adjacent sides of a rhombus have lengths (2x+11)cm and (4x+5)cm. Find x and hence find the perimeter of the rhombus
Answer by flame8855(424)

2x+11 = 4x +5
2x = 6
x = 3
side = 17
perimeter = 17*4 = 68 cm